Given that x=2cosθ and y=2sinθ evaluate x2+y2


1
Expert's answer
2021-10-27T11:25:38-0400

Solution:

"x=2\cosθ , y=2\sinθ"

On squaring both sides of each equation:

"x^2=2^2\cos^2θ ,y^2=2^2\sin^2θ
\\ \Rightarrow x^2=4\cos^2θ ,y^2=4\sin^2θ"

Now,

"x^2+y^2=4\cos^2θ +4\sin^2θ
\\=4(\cos^2θ +\sin^2θ)
\\=4(1) \ \ [\because \cos^2A +\sin^2A=1] 
 \ \ \\=4"